# GigaSpeech 2: An Evolving, Large-Scale and Multi-domain ASR Corpus for Low-Resource Languages with Automated Crawling, Transcription and Refinement

### v1

@@ 4.1 ASR Model Training on GigaSpeech 2
Our ASR systems are constructed by Zipformer Transducer [16]. Two Zipformer [48] variants, namely Zipformer-M and Zipformer-L, are employed for each NST iteration. Specific configurations are listed in Appendix C.1. During Noisy Student Training, SpecAugment [31] is used as input noise, and Bypass [48] and feature mask [48] are used as model noise.
Table 2: Comparison of ASR performance with different NST iterations on various evaluation sets, including GigaSpeech 2 DEV and TEST, Common Voice 17.0 TEST, and FLEURS TEST. Detailing training set size (# Hours), model size (# Params), Character Error Rate (CER) for Thai, and Word Error Rate (WER) for Indonesian and Vietnamese.
@@
Table 2 presents the ASR results across different NST iterations on three evaluation sets, including the development and test sets from GigaSpeech 2 and the Common Voice 17.0 and FLEURS test set. Each iteration involves distinct modifications aimed at refining high-quality transcriptions. A subset of automatic transcriptions generated by Whisper large-v3 is used to train the initial teacher model (Iteration 1). The teacher model then filters the training utterances by applying a CER/WER threshold, using the original labels as references and the new labels generated by the teacher as the hypothesis. The student model is trained on this filtered set with noise injected (Iteration 2). The student model is then used as the teacher to generate new labels on a larger subset of raw automatic transcriptions, applying the same filter to refine the training data. This refined data is used to train the student model with noise injected (Iteration 3). The process repeats in subsequent iterations, and the model size is scaled up to a larger version in the final iteration (Iteration 3 of Indonesian and Vietnamese, Iteration 4 of Thai).
@@
1) Across all three languages (Thai, Indonesian, and Vietnamese), iteratively scaling the training data size, adding noise, and filtering labels lead to consistent improvements in the WER performance on the evaluation sets until the final iteration. This indicates that the iterative approach of refining and scaling the training data is effective in enhancing the accuracy of the raw transcriptions.
2) The Thai language achieves the absolute lowest error rates consistently across iterations from Iteration 1 to 4, indicating the effectiveness of the NST approach for this particular language. The best NST model outperforms the standard transcription model data by WER reductions of 1.69%, 2.64%, 4.73%, and 3.79% absolute (13.92%, 17.48%, 53.27%, and 26.45% relative) respectively (Iteration 4 vs. Iteration 1).
Additional ablation studies on our modified NST in Appendix D Table 8 demonstrate the effectiveness of relabeling and discuss the detriment of enlarging noise when scaling the training data.
4.2 Comparison to Existing ASR Systems
To demonstrate the efficacy of our ASR models trained on Gigaspeech 2, several mainstream and competitive ASR systems, including Whisper [37] from OpenAI, MMS [34] from Meta, and commercial services from Azure and Google, are used as benchmarks.
Whisper: Our work builds upon Whisper [37], a suite of large-scale, multitask, and multilingual speech models developed by OpenAI. It leverages the encoder-decoder Transformer architecture [45], with model sizes ranging from 39 million parameters (tiny) to 1.55 billion parameters (large). Additionally, Whisper offers variants spanning from an English-only version to a multilingual model capable of handling 99 languages. To conduct a comprehensive evaluation, we test three configurations: Whisper base, Whisper large-v2, and Whisper large-v3 models.
MMS: The Massively Multilingual Speech (MMS) [34] project leverages self-supervised learning (SSL) techniques and a novel dataset to expand the language coverage of speech technology significantly. The core components include pre-trained wav2vec 2.0 [2] models for 1,406 languages, a single multilingual ASR model supporting 1,107 languages, speech synthesis models for the same set of languages, and a language identification model capable of recognizing 4,017 languages. In this study, we employ the MMS L1107 configuration.
Google USM: The Universal Speech Model (USM) [50] is introduced as a single, large-scale model that excels in ASR across over 100 languages. This achievement is made possible by pre-training the model’s encoder on a vast, unlabeled multilingual dataset of 12 million hours, covering more than 300 languages, followed by fine-tuning on a smaller labeled dataset. To conduct a thorough comparison, we utilize their Chirp Speech-to-Text v2 model for performance evaluation.
Table 3: Comparison of ASR results for models trained on GigaSpeech 2 with open-source multilingual ASR models and commercial ASR services, evaluated on test sets from GigaSpeech 2, Common Voice 17.0, and FLEURS. The evaluation metrics are Character Error Rate (CER) for Thai and Word Error Rate (WER) for both Indonesian and Vietnamese. $\dagger$" denotes commercial services.
@@ 4.2 Comparison to Existing ASR Systems
We compare the performance of our proposed approach trained on Gigaspeech 2 against these above-mentioned ASR models, including Whisper (base, large-v2, and large-v3), MMS L1107, Azure Speech CLI 1.37.0 and Google USM Chirp v2  [^6] Abnormal high deletion rates with Google USM in Thai are observed in our repeated testing. , across three languages: Thai, Indonesian, and Vietnamese. The ASR performance is evaluated regarding character error rate (CER) or word error rate (WER) on three distinct test sets from GigaSpeech 2, Common Voice 17.0, and FLEURS. According to the results shown in Table 3, there are several intriguing findings:
1) For the Thai language, our ASR model trained on Gigaspeech 2 (Table 3, Thai, Row 7) outperforms all competitors, including commercial services from Azure and Google USM, securing the top rank across all three test sets among the seven models. It outperforms Whisper large-v3 by WER reductions of 7.98%, 1.87%, and 1.01% absolute (39.04%, 31.06%, and 8.74% relative) (Table 3, Thai, Row 7 vs. Row 1). Remarkably, our model achieves such impressive performance with nearly one-tenth of the parameters compared to Whisper large-v3 (151.9 M vs. 1542.0 M).
2) For the Indonesian and Vietnamese languages, our system demonstrates competitive performance compared to existing baseline models. This highlights the efficacy of our pipeline in delivering high-quality results with a lightweight model. Specifically, on the GigaSpeech 2 test set in the Indonesian language, our system (Table 3, Indonesian, Row 7) outperforms all baseline models, attaining the best performance. Compared to Whisper large-v3, the model trained on Indonesian achieves an absolute WER reduction of 5.11%, corresponding to a relative reduction of 25.51% (Table 3, Indonesian, Row 7 vs. Row 1). Similarly, the model trained on Vietnamese achieves an absolute WER reduction of 5.11% compared to the Whisper large-v3 model. (Table 3, Vietnamese, Row 7 vs. Row 1).
3) Our model exhibits degraded performance compared to commercial ASR systems on the Common Voice and FLEURS test sets in Indonesian and Vietnamese, which can be attributed to the domain mismatch. Contrastively, we observe a performance leap after adding Common Voice and FLEURS training data into GigaSpeech 2 (Table 3, Indonesian & Vietnamese, Row 7 vs. Row 8).
Despite this substantial disparity in training data size, our method achieves the best performance for the Thai language domain and delivers comparable results to commercial models for Indonesian and Vietnamese. This remarkable accomplishment highlights the efficacy of our approach in leveraging limited, free, open-source, unlabeled data to train highly competitive speech recognition models. It showcases a promising path towards developing high-quality speech recognition systems without the need for extensive, proprietary datasets, thereby reducing the barrier to entry and enabling wider accessibility.
4.3 Comparison to the YODAS corpus
Table 4: Comparison of ASR results for models trained on YODAS and GigaSpeech 2, evaluated on test sets from GigaSpeech 2, Common Voice 17.0, and FLEURS. The evaluation metrics are Character Error Rate (CER) for Thai and Word Error Rate (WER) for both Indonesian and Vietnamese.
@@ 4.3 Comparison to the YODAS corpus
Table 4 compares ASR performance across different models trained on YODAS [29] and GigaSpeech 2 datasets evaluated on various test sets. Note that YODAS Thai automatic is not included because of insufficient data (only 1 hour). Analyzing the trends, the following observations can be concluded:
1) The models trained on GigaSpeech 2 refined yield significantly better results than models trained on the YODAS datasets in the GigaSpeech 2 test set for all three languages.
2) For Thai and Vietnamese, the models trained on GigaSpeech 2 refined consistently outperform those trained on the YODAS manual datasets across all evaluation sets. For Thai, the best model outperforms the YODAS manual with relative WER reductions of 56.68%, 62.13%, and 34.57% (Table 4, Thai, Row 3 vs. Row 2). For Vietnamese, it achieves relative WER reductions of 68.48%, 55.71%, and 60.47% relative (Table 4, Vietnamese, Row 4 vs. Row 3).
3) The YODAS manual may suffer from overfitting or noisy data issues due to simplistic filtering rules, leading to inconsistent performance in Indonesian.
4) Purely automatic generation of YODAS tends to degrade performance, as observed for Vietnamese and Indonesian, likely due to the inherent noise and errors in the automatically generated subtitles.
4.4 Training ASR Models within ESPNet and Icefall on GigaSpeech 2
Icefall: The neural Transducer [16] architecture is adopted, with Zipformer-L as the encoder and the pruned RNN-T loss [26] as the object function. 2000-class Byte Pair Encoding (BPE) [39] word pieces are used. More details are provided in Appendix C.1.
ESPNet: The Conformer [17] CTC/AED [25] system is adopted from ESPNet [46], with Conformer-L as the encoder and a combination of the localized sensitivity of convolutional neural networks and the long-range modeling capabilities of Transformers [45]. 2000-class BPE word pieces are used. More details can be found in Appendix C.2.
Table 5: Comparison of ASR models trained on GigaSpeech 2 with Icefall and ESPNet toolkits, evaluated on GigaSpeech 2 TEST set. The evaluation metrics are Character Error Rate (CER) for Thai (th) and Word Error Rate (WER) for both Indonesian (id) and Vietnamese (vi).
@@ 4.4 Training ASR Models within ESPNet and Icefall on GigaSpeech 2
ESPNet Conformer/Transformer CTC/AED 111.8 13.70 15.50 15.60
@@
5 Limitation and Future Work
Due to time constraints, we only tested 3-4 iterations of the proposed NST model. We are optimistic that more iterations will yield even better results. The Indonesian and Vietnamese datasets are not as extensive as the Thai datasets, so we plan to add more data for these two languages. We will also expand our low-resource language family in our future investigation. In addition, we did not perform language model fusion to further boost performance since there is a lack of high-quality and in-domain text data for low-resource languages. To resolve potential legal risks, our dataset follows the same license as GigaSpeech [6], allowing only non-commercial research and educational use.
@@ 6 Conclusion
This paper introduces a new multilingual speech dataset, GigaSpeech 2, and a novel automated pipeline to boost speech recognition performance using in-the-wild audio-only data. GigaSpeech 2 aims to address the scarcity of labeled training data on low-resource languages by developing this large-scale, multi-domain, and multilingual corpus. Extensive experiments are conducted to validate the efficacy of our newly introduced corpus. The ASR models trained in three languages, which are Thai, Indonesian, and Vietnamese within GigaSpeech 2, demonstrate superior and impressive performance compared to various powerful ASR models, including Whisper large v2/v3 from OpenAI, MMS from Meta, and even commercial services from Google and Azure. The related resources, including the training corpus, curated test sets, automated pipeline, and recipes, will be released  [^7] https://github.com/SpeechColab/GigaSpeech2 later to facilitate research in this direction. In the future, we are eager to extend our paradigm to more low-resource languages and devoted to breaking down the language barrier.
Acknowledgments and Disclosure of Funding
We thank DataOcean AI for their manual annotation of the evaluation set.
Appendix A Algorithm of Iterative Label Refinement
Algo. 1 illustrates the workflow of our proposed iterative label refinement.
Output: Refined-label set $\mathcal{R}$
$\mathcal{R}\leftarrow\mathcal{P}_{1}$;
Train teacher model $\mathcal{M}_{1}$ on $\mathcal{R}$ with noise;
for $i\leftarrow 1$ to $n$ do
       $\mathcal{R}\leftarrow\varnothing$;
       if $i$ == $1$  then
             // Filter $\mathcal{P}_{i}$ by teacher model $\mathcal{M}_{i}$ with CER $\leq\tau$
      else
             for $j\leftarrow 1$ to $i$ do
                   $\mathcal{R}\leftarrow\mathcal{R}\cup\mathcal{R}_{tmp}$;
             end for
       end if
end for
return $\mathcal{R}$;
Algorithm 1 Iterative Label Refinement
Appendix B Detailed Analysis of GigaSpeech 2
Detailed analysis of the GigaSpeech 2 dataset is presented through the following visualizations.
B.1 Domain Distribution of Manual Evaluation Sets
The domain distribution of the manual evaluation sets is shown in Fig. 2.
(a) Thai
(b) Indonesian
(c) Vietnamese

### v2

@@ 4.1 ASR Model Training on GigaSpeech 2
Our ASR systems are built on Zipformer Transducer Graves et al. (2013). Two Zipformer Yao et al. (2024) variants, namely Zipformer-M and Zipformer-L, are employed for each NST iteration. Specific configurations are provided in Appendix B.1. During Noisy Student Training, SpecAugment Park et al. (2019) is used as input noise while Bypass Yao et al. (2024) and feature mask Yao et al. (2024) are used as model noise.
Table 2: Comparison of ASR performance across different NST iterations on various evaluation sets, including GigaSpeech 2 DEV and TEST, Common Voice 17.0 TEST, and FLEURS TEST. Reported details include training set size (#Hours), BPE vocabulary size (#Vocab), model size (#Params), CER for Thai, and WER for Indonesian and Vietnamese.
@@
Table 2 presents the ASR results across different NST iterations on three evaluation sets, including the development and test sets from GigaSpeech 2 and the Common Voice 17.0 and FLEURS test set. Each iteration involves distinct modifications aimed at refining high-quality transcriptions. A subset of automatic transcriptions generated by Whisper large-v3 is used to train the initial teacher model (Iter. 1). The teacher model then filters the training utterances by applying a CER/WER threshold, using the original labels as references and the new labels generated by the teacher as the hypothesis. The student model is trained on this filtered set with noise injected (Iter. 2). The student model is then used as the teacher to generate new labels on a larger subset of raw automatic transcriptions, applying the same filter to refine the training data. This refined data is used to train the student model with noise injected (Iter. 3). The process repeats in subsequent iterations, and the model size is scaled up to a larger version in the final iteration (Iter. 3 of Indonesian & Vietnamese, Iter. 4 of Thai).
@@
1) Across all three languages, iteratively scaling the training data size, adding noise, and filtering labels lead to consistent improvements in the WER performance on the evaluation sets until the final iteration. This indicates that the iterative approach of refining and scaling the training data is effective in enhancing the accuracy of the raw transcriptions.
2) The system trained on Thai consistently achieves the absolute lowest error rates consistently across iterations from 1 to 4, indicating the effectiveness of the NST approach for this particular language. The best NST model outperforms the standard transcription model data by WER reductions of 1.69%, 2.64%, 4.73%, and 3.79% absolute (13.92%, 17.48%, 53.27%, and 26.45% relative) respectively (Iter. 4 vs. 1).
Additional ablation studies on our modified NST in Appendix C Table 8 demonstrate the effectiveness of relabeling and discuss the detriment of enlarging noise when scaling the training data.
Table 3: Comparison of ASR results for models trained on GigaSpeech 2 with open-source multilingual ASR models and commercial ASR services, evaluated on three test sets from GigaSpeech 2, Common Voice 17.0, and FLEURS. The evaluation metrics include CER for Thai and WER for both Indonesian and Vietnamese. $\dagger$" denotes commercial services.
@@
Table 4: Comparison of ASR results for models trained on YODAS and GigaSpeech 2, evaluated on test sets from GigaSpeech 2, Common Voice 17.0, and FLEURS. The evaluation metrics include CER for Thai and WER for both Indonesian and Vietnamese.
@@
Table 5: Comparison of ASR models trained on GigaSpeech 2 with Icefall and ESPnet toolkits, evaluated on GigaSpeech 2 TEST set. The evaluation metrics include CER for Thai (th) and WER for both Indonesian (id) and Vietnamese (vi).
@@
ESPnet Conformer/Transformer CTC/AED 111.8 13.70 15.50 14.60
4.2 Comparison to Existing ASR Systems
To demonstrate the efficacy of our ASR models trained on GigaSpeech 2, several mainstream and competitive ASR systems, including Whisper Radford et al. (2023) from OpenAI, MMS Pratap et al. (2024) from Meta, and commercial services from Azure and Google, are used as benchmarks.
Whisper: Our work builds upon Whisper Radford et al. (2023), a suite of large-scale, multitask, and multilingual speech models developed by OpenAI. It leverages the encoder-decoder Transformer architecture Vaswani et al. (2017), with model sizes ranging from 39 million parameters (tiny) to 1.55 billion parameters (large). Additionally, Whisper offers variants spanning from an English-only version to a multilingual model capable of handling 99 languages. To conduct a comprehensive evaluation, we test three variants: Whisper base, Whisper large-v2, and Whisper large-v3 models.
MMS: The Massively Multilingual Speech (MMS) Pratap et al. (2024) project leverages self-supervised learning (SSL) techniques and a novel dataset to expand the language coverage of speech technology significantly. The core components include pre-trained wav2vec 2.0 Baevski et al. (2020) models for 1,406 languages, a single multilingual ASR model supporting 1,107 languages, speech synthesis models for the same set of languages, and a language identification model capable of recognizing 4,017 languages. In this study, we employ the MMS L1107 configuration.
Google USM: The Universal Speech Model (USM) Zhang et al. (2023) is introduced as a single, large-scale model that excels in ASR across over 100 languages. This achievement is made possible by pre-training the model’s encoder on a vast, unlabeled multilingual dataset of 12 million hours, covering more than 300 languages, followed by fine-tuning on a smaller labeled dataset. To conduct a thorough comparison, we utilize their Chirp Speech-to-Text v2 model for performance evaluation.
We compare the performance of our proposed approach trained on GigaSpeech 2 against these above-mentioned ASR models, including Whisper (base, large-v2, and large-v3), MMS L1107, Azure Speech CLI 1.37.0 and Google USM Chirp v2666Abnormal high deletion rates with Google USM in Thai are observed in our repeated testing., across three languages: Thai, Indonesian, and Vietnamese. The ASR performance is evaluated regarding character error rate (CER) or word error rate (WER) on three distinct test sets from GigaSpeech 2, Common Voice 17.0, and FLEURS. According to the results shown in Table 3, there are several intriguing findings:
1) For the Thai language, our ASR model trained on GigaSpeech 2 (Table 3, Thai, Row 7) outperforms all competitors, including commercial services from Azure and Google, securing the top rank across all three test sets among the seven models. It outperforms Whisper large-v3 by relative WER reductions of 39.04%, 31.06%, and 8.74% (Table 3, Thai, Row 7 vs. 1). Remarkably, our model achieves such impressive performance with nearly one-tenth of the parameters compared to Whisper large-v3 (151.9 M vs. 1542 M).
2) For the Indonesian and Vietnamese languages, our system demonstrates competitive performance compared to existing baseline models. This highlights the efficacy of our pipeline in delivering high-quality results with a lightweight model. Specifically, on the GigaSpeech 2 test set in the Indonesian language, our system (Table 3, Indonesian, Row 7) outperforms all baseline models, attaining the best performance. Compared to Whisper large-v3, the model trained on Indonesian achieves an absolute WER reduction of 5.11%, corresponding to a relative reduction of 25.51% (Table 3, Indonesian, Row 7 vs. 1). Similarly, the model trained on Vietnamese achieves an absolute WER reduction of 5.11%, corresponding to a relative reduction of 28.48% (Table 3, Vietnamese, Row 7 vs. 1).
3) Our model exhibits degraded performance compared to commercial ASR systems on the Common Voice and FLEURS test sets in Indonesian and Vietnamese, which can be attributed to the domain mismatch [^7] Unlike GigaSpeech 2, which contains noisy, reverberant spontaneous speech, Common Voice and FLEURS comprise clean, read speech with text from written materials.. Contrastively, we observe a performance leap after adding Common Voice and FLEURS training data into GigaSpeech 2 (Table 3, Indonesian & Vietnamese, Row 7 vs. 8).
Although our training data size is smaller than that of industrial-scale models, our method achieves the best performance for the Thai language domain and delivers comparable results to commercial models for Indonesian and Vietnamese. This remarkable accomplishment highlights the efficacy of our approach in leveraging limited, free, open-source, unlabeled data to train highly competitive speech recognition models. It showcases a promising path towards developing high-quality speech recognition systems without the need for extensive, proprietary datasets, thereby reducing the barrier to entry and enabling wider accessibility.
4.3 Comparison to the YODAS Corpus
Table 4 compares ASR performance across different models trained on YODAS Li et al. (2023) and GigaSpeech 2 datasets evaluated on multiple test sets. Note that YODAS Thai automatic is not included due to insufficient data (only 1 hour). Despite variations in overall data volume, several general conclusions can be drawn from trend analysis:
1) The models trained on GigaSpeech 2 refined yield generally superior results compared to those trained on YODAS datasets for all three languages.
2) The YODAS manual may suffer from overfitting or noisy data issues due to simplistic filtering rules, leading to inconsistent performance in Indonesian (Table 4, Indonesian, Row 1 & 3).
3) Purely automatic generation of YODAS tends to degrade performance, as observed for Vietnamese (Table 4, Vietnamese, Row 1 vs. 2) and Indonesian (Table 4, Indonesian, Row 1 vs. 2), likely due to the inherent noise and errors in the automatically generated subtitles.
4.4 Training ASR Models within ESPnet and icefall on GigaSpeech 2
Icefall: We adopt the neural Transducer Graves et al. (2013) architecture, using Zipformer-L as the encoder, the pruned RNN-T loss Kuang et al. (2022) as the object function, and 2000-class Byte Pair Encoding (BPE) Sennrich et al. (2016) word pieces. More details are provided in Appendix B.1.
ESPnet: We employ Conformer Gulati et al. (2020) CTC/AED Kim et al. (2017) system from ESPnet Watanabe et al. (2018), using Conformer-L as the encoder and 2000-class BPE word pieces. This model combines the localized sensitivity of convolutional neural networks with the long-range modeling capabilities of Transformers Vaswani et al. (2017). Details are available in Appendix B.2.
@@ 5 Conclusion
This paper introduces a new multilingual speech dataset, GigaSpeech 2, and a novel automated pipeline to boost speech recognition performance using in-the-wild audio-only data. GigaSpeech 2 aims to address the scarcity of labeled training data on low-resource languages by developing this large-scale, multi-domain, and multilingual corpus. Extensive experiments are conducted to validate the efficacy of our newly introduced corpus. The ASR models trained in three languages, which are Thai, Indonesian, and Vietnamese within GigaSpeech 2, demonstrate superior and impressive performance compared to various powerful ASR models, including Whisper large v2/v3 from OpenAI, MMS from Meta, and even commercial services from Google and Azure. The related resources, including the corpus with curated test sets [^8] https://huggingface.co/datasets/speechcolab/gigaspeech2, automated pipeline [^9] https://github.com/SpeechColab/GigaSpeech2, and recipes [^10] https://github.com/k2-fsa/icefall, are released to facilitate research in this direction. In the future, we are eager to extend our paradigm to more low-resource languages and are devoted to breaking down the language barrier.
Limitations
In this paper, we propose GigaSpeech 2, a large-scale, multi-domain, multilingual speech recognition corpus, and a novel automated pipeline to boost speech recognition performance using in-the-wild audio-only data. We only conducted 3-4 iterations of the proposed NST method in our experiments, and we are optimistic that more iterations on large data will yield even better results. Moreover, we are actively extending our language coverage by incorporating additional languages, including Malay, Korean, Minnan, and Arabic. We will also expand our low-resource language family in our future investigation.
Ethics Statement
All collected audio is sourced from materials released under a Creative Commons license. Personally identifiable information has been anonymized using rule-based scripts to remove identifiable content from the data. All annotators are compensated fairly by a professional data annotation company. Our dataset adopts the same terms as GigaSpeech Chen et al. (2021) to resolve potential legal risks, restricting use to non-commercial research and educational purposes only. We are committed to ongoing maintenance of the dataset to address any potential risks in the future.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. U23B2018 and No. 62206171), Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0102 and Yangtze River Delta Science and Technology Innovation Community Joint Research Project (2024CSJGG01100). We gratefully acknowledge the support of DataOcean AI for manually annotating the evaluation sets.
Appendix A Detailed Analysis of GigaSpeech 2
A.1 Manual Transcription Quality Assurance
The manual transcription process, carried out by a professional data annotation company, includes rigorous manual quality checks and secondary inspections to ensure that timestamp accuracy and transcription correctness exceed 97%. All manually transcribed results undergo a 100% manual quality inspection, where both timestamps and transcription accuracy are thoroughly checked. Any data that fails to meet the required standards is sent back for correction. Subsequently, 30% of each inspector’s reviewed data is re-evaluated. If this recheck confirms over 97% accuracy, the data passes; otherwise, the entire dataset inspected by that quality inspector is returned for full correction. For timestamp accuracy, an audio snippet tool is used to ensure that timestamps do not overlap with the waveform. If any timestamp does fall on the waveform, a manual inspection is conducted to confirm whether it corresponds to speech.
A.2 Domain Distribution of Manual Evaluation Sets
The domain distribution of the manual evaluation sets is shown in Fig. 2. The domains are identified based on a predefined set of categories. Each sample is manually annotated at the individual video level, considering both the topic type and content format.
(a) th
(b) id
(c) vi